OpenStudy (anonymous):

Lizbeth deposited $6200 into a savings account for which interest is compounded monthly at a rate of 2.37%. How much interest will she earn after 8 years?

OpenStudy (tkhunny):

Have you considered the accumulation formula? \(6200\cdot\left(1 + \dfrac{0.0237}{12}\right)^{12\cdot 8}\)